OpenStudy (anonymous):

Can somebody please help me with these last 4! Im stuck, I've tried over and over :( Given the function f(x) = The quantity of 3x minus 4, divided by 5, which of the below expressions is correct? f−1(x) = The quantity of 5x plus 4, divided by 3. f−1(x) = The quantity of 5x minus 4, divided by 3. f−1(x) = The quantity of negative 3x minus 4, divided by 5. f−1(x) = The quantity of 4 minus 3x, divided by 5.

OpenStudy (solomonzelman):

kinda too many questions per thread. I will help you with the first one here, if you do not mind.

OpenStudy (anonymous):

Thats fine, I can post in a seperate thread if needed :) thank you so much

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x) = \frac{ 3x - 4}{5} }\)

OpenStudy (solomonzelman):

first, replace f(x) with "y" (for convenience) then change x and y with each other (put x instead of y, and vice versa)

OpenStudy (solomonzelman):

and then solve for y in your new equation. (after you find the y, denote this y as \(f^{-1}(x)\) )

OpenStudy (solomonzelman):

can you do the first step 1) replace \(f(x)\) with y

OpenStudy (anonymous):

So now it is x= 3y-4 / 5

OpenStudy (solomonzelman):

yes x=(3y-4)/5

OpenStudy (solomonzelman):

then you need to isolate the y

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

Do i multiply by 5?

OpenStudy (solomonzelman):

yes, that is good!

OpenStudy (solomonzelman):

multiply times 5 on both sides

OpenStudy (anonymous):

So it would look like (5) x = 3y-4 / 5 (5)

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ \displaystyle x=\frac{ 3y-4 }{5} }\) \(\large\color{black}{ \displaystyle x\color{blue}{\times 5}=\frac{ 3y-4 }{5} \color{blue}{\times 5} }\)

OpenStudy (solomonzelman):

continue please:)

OpenStudy (anonymous):

Now im stuck at this part :(

OpenStudy (solomonzelman):

the 5's on the right cancel, and the left side is just 5x \(\large\color{black}{ \displaystyle x \color{blue}{\times 5}=\frac{ 3y-4 }{\cancel{5}}\color{blue}{\times \cancel{5}} }\) \(\large\color{black}{ \displaystyle 5x = 3y-4 }\)

OpenStudy (solomonzelman):

making sense?

OpenStudy (anonymous):

Yes, now that makes sense

OpenStudy (solomonzelman):

ok, now isolate the y

OpenStudy (solomonzelman):

(or, "solve for y" -- same thing)

OpenStudy (anonymous):

Would I start by adding the 4 to 5x?

OpenStudy (solomonzelman):

yes, good!

OpenStudy (solomonzelman):

what is your equation now?

OpenStudy (anonymous):

9x = 3y?

OpenStudy (solomonzelman):

oh, 5x and 4 aren't like terms

OpenStudy (anonymous):

I mean 9 = 3y :)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 5x = 3y-4 }\) you are adding 4 to both sides. \(\large\color{black}{ \displaystyle 5x\color{red}{+4} = 3y-4 \color{red}{+4} }\) what happens to the right side? the left side side remains the way it is right now, because \(5x\) and \(4\) can't be added together.

OpenStudy (anonymous):

the right side cancels out so we're left with 5 x +4 /3 would be the final asnwer?

OpenStudy (solomonzelman):

that would be the final answer.

OpenStudy (solomonzelman):

nicely done

OpenStudy (anonymous):

Thank you so much! :) I think I understand this mess a little better now lol

OpenStudy (solomonzelman):

\(\normalsize\color{ slate }{\Huge\displaystyle {\bbox[5pt, lightcyan ,border:2px solid black ]{ {\rm f}^{-1}(x)=\frac{5x+4}{3} }}}\)

OpenStudy (solomonzelman):

basically, all you need to do in general is: 1) replace f(x) with y. 2) change x and y (replace them with each other- x for y, and y for x) 3) {solve for/ isolate the} y (this is the difficult part) 4) after you found y, denote this y as an inverse function, because this y is indeed the inverse function. So instead of y, write \({\rm f}^{-1}(x)\).

OpenStudy (solomonzelman):

glad that it helped. I have to depart for some time.... see you.

OpenStudy (anonymous):

I think I got it, thank you again :)